### v1

# Racka: Efficient Hungarian LLM Adaptation on Academic Infrastructure

Abstract
We present Racka, a lightweight, continually pretrained large language model designed to bridge the resource gap between Hungarian and high‑resource languages such as English and German. Racka employs parameter‑efficient continual pretraining via Low‑Rank Adaptation (LoRA) on a Qwen‑3 4B backbone, making the recipe practical on A100 (40GB)‑based HPC clusters with low inter‑node bandwidth. To better match the training distribution, we replace and adapt the tokenizer, achieving substantially improved tokenization fertility for Hungarian while maintaining competitive performance in English and German. The model is trained on 160B subword tokens drawn from a mixture of internet and high‑quality curated sources, with a composition of 44% Hungarian, 24% English, 21% German, and 11% code. This data mix is chosen to mitigate catastrophic forgetting and preserve high‑resource language capabilities during continual pretraining. Our preliminary results indicate modest but stable results in language adaptation.
Keywords: Large Language Model, Continual Pretraining, Language adaptation, Tokenizer Adaptation
1 Introduction
1.1 The Challenge of Linguistic Digital Sovereignty for Medium-Resource Languages
The proliferation of Large Language Models (LLMs) has marked a transformative era in artificial intelligence, yet this progress has been unevenly distributed across the world’s languages (ali2024surveylargelanguagemodels). The current landscape is dominated by models pretrained on vast corpora composed predominantly of English and a few other high-resource languages, creating a significant performance and resource disparity for less-resourced linguistic communities (zhong2025opportunitieschallengeslargelanguage). For medium-resource languages such as Hungarian, a Finno-Ugric language characterized by its agglutinative nature and rich morphology, this gap is particularly pronounced. Off-the-shelf multilingual models often exhibit suboptimal performance due to insufficient representation in training data and tokenizers that are ill-suited to their linguistic structure. This is particularly the case for open-source models, which visibly struggle with Hungarian grammar.
This disparity underscores a critical challenge to linguistic digital sovereignty: the capacity for a linguistic community to develop, deploy, and benefit from AI technologies that are not only proficient in its language but are also attuned to its unique cultural and contextual nuances. Achieving this requires the creation of models trained on nationally relevant data and, crucially, developed using methodologies that are accessible to local research communities, who may not have access to the hyperscale computing infrastructure available to large industrial labs, making researchers rely on shared HPC infrastructure.
1.2 Core Contributions
In response to these challenges, we introduce Racka, a 4-billion-parameter language model for Hungarian adapted trough a pragmatic and resource-efficient approach from a Qwen3-4B reasoning model. The model’s name pays homage to the Hortobágy Racka, an indigenous Hungarian sheep breed and llama alternative. Racka was trained using publicly accessible infrastructure—Hungary’s Komondor supercomputer—rather than state-of-the-art proprietary hardware.
This paper presents a comprehensive recipe for language adaptation that is both effective and practical for research communities with access to national HPC clusters. Our primary contributions are threefold:
  1. 1.
    A Resource-Efficient Training Recipe: We detail a practical methodology for the continual pretraining of a 4B parameter model on an HPC cluster featuring NVIDIA A100 (40GB) GPUs. We provide justification for our selection of Distributed Data Parallel (DDP) over the more recent Fully Sharded Data Parallel (FSDP) paradigm, demonstrating that this choice is a deliberate optimization for our specific hardware environment.
  2. 2.
    Optimized Tokenization for Hungarian: We underscore the critical role of tokenizer adaptation for morphologically rich, agglutinative languages. We document the design, training, and evaluation of a new tokenizer that dramatically improves tokenization efficiency (i.e., subword fertility) for Hungarian, while carefully preserving the performance of the base model on high-resource control languages like English and German.
  3. 3.
    Curation of a Large-Scale Multilingual Corpus: We describe the assembly and curation of a 160 billion subword token dataset representing state-of-the-art size in training a Hungarian-centric language model. We justify its strategic composition (44% Hungarian, 43% high-resource languages, and 11% code) as a pragmatic approach to mitigate catastrophic forgetting during the continual pretraining process.
2 Related Work
2.1 Paradigms of Language Model Adaptation
Adapting existing foundation models to new languages, domains, or tasks is a central challenge in modern NLP. The prohibitively high cost of training LLMs from scratch has spurred research into more efficient adaptation strategies. (csaki2023efficientlyadaptingpretrainedlanguage; khade2024challengesadaptingmultilingualllms) As an industry standard practice for language adaptation, continual pretraining (CPT) is widely adopted. This method focuses on further self-supervised training on large, unlabeled corpora to expand a model’s foundational knowledge. (vanroy2024geitje7bultraconversational; laion2024leolm; ke2023continual)
The work on Racka falls squarely within the CPT paradigm, specifically targeting what the literature defines as “Language Expansion” and “Domain Adaptation”. This framing positions our effort not as the creation of a final, task-specific model, but as the development of a more capable foundational model for Hungarian, which can subsequently be adapted for downstream tasks. A primary obstacle in any continual learning scenario is catastrophic forgetting, where a model’s performance on previously learned tasks or languages degrades significantly as it learns new information (gupta2023continual). A simple yet effective strategy to mitigate this is data replay, which involves mixing data from the original training distribution with the new data. Our multilingual data composition, which retains a significant portion of English and German text, is a direct application of this principle. Other alternatives include careful train scheduling, intense regularizations, or the analysis of task-vector shifts. We refrain from these methods in favor of data replay, as it directly controls our training objective and does not require extensive experimentation (li2024examiningforgettingcontinualpretraining). Previous Hungarian language adaptation work also applied this approach successfully (szentmihalyi2024pretraining; csaki2024sambalingoteachinglargelanguage; gyozo2025pulillumix).
2.2 Parameter-Efficient Methods for Continual Pretraining
The computational burden of updating the billions of parameters in an LLM has led to the development of Parameter-Efficient Fine-Tuning (PEFT) methods. These techniques aim to adapt models by training only a small fraction of their total parameters, drastically reducing memory and compute requirements. Among the most widely adopted PEFT methods is Low-Rank Adaptation (LoRA). (hu2021loralowrankadaptationlarge) LoRA operates on the hypothesis that the change in model weights during adaptation has a low intrinsic rank. This aligns well with our aim to train on limited-memory GPUs with low throughput internode connections. By keeping the base model frozen, it is shown to inherently aid in overcoming catastrophic forgetting of the base model’s knowledge thus making them ideal for CPT (li2024examiningforgettingcontinualpretraining).
2.3 The Landscape of Hungarian Language Models
The development of Racka builds upon a growing body of work dedicated to creating capable language models for Hungarian. These efforts span both academic and industrial research and provide essential context for our contributions.
  • NYTK Models: The Hungarian Research Centre for Linguistics (NYTK) has been a key contributor, developing a series of models under the “PULI” family. These include models in the range of 7-8B parameters, both monolingual and continually pretrained. Latest models include instruction-tuned and multilingual (English, Chinese) models. (yang-puli-gptrio; gyozo2025pulillumix; yang-llumix-llama)
  • OTP Bank Models: A major industrial and government supported initiative resulted in the OTP 1.5B and OTP 13B bilingual models, which were trained on an extensive dataset in collaboration with SambaNova Systems. These models demonstrated the feasibility of developing high-performance models for Hungarian at a large scale. (szentmihalyi2024pretraining)
  • International Initiatives: The ecosystem also includes models adapted from multilingual open-source foundations. SambaLingo-Hungarian-Base, for instance, adapted a Llama-2 7B model by training on 59 billion tokens from the Hungarian portion of the Cultura-X dataset. (csaki2024sambalingoteachinglargelanguage) Similarly, projects like OpenEuroLLM have fine-tuned models like Gemma for improved Hungarian performance. Other European initiatives such as EuroLLM and Salamandra include Hungarian in their supported languages. (openeurollm2025hungarian; eurollm2025; gonzalezagirre2025salamandratechnicalreport)
This existing landscape highlights a clear and valuable niche for the Racka project. While previous efforts have focused on larger models (OTP-13B), monolingual training (PULI-GPT-3SX), or adaptation on smaller datasets (SambaLingo), Racka’s contribution is its unique combination of a lightweight (4B) backbone, and a continual pretraining on an exceptionally large multilingual corpus (160B tokens), and a resource-efficient methodology explicitly designed for and validated on publicly accessible national HPC infrastructure. One key insight tha can be accrued from the papers above is that for smaller languages, continual pretraining of already capable models results in better performning LLMs than training new models from scratch.
3 Training Data Curation
3.1 Corpus Design and Composition
The foundation of any successful language model is the quality and scale of its training data. For Racka, we curated a comprehensive multilingual corpus totaling approximately 160 billion subword tokens with a dataset size of $639$ GB. The composition of this corpus was a strategic design choice aimed at achieving robust language adaptation for Hungarian while simultaneously mitigating the effects of catastrophic forgetting on the backbone model’s existing capabilities. The final data mixture, detailed in Table 1, allocates the majority of the data to our target language, Hungarian, while retaining substantial portions of high-quality English, German, and source code data. The inclusion of English and German serves as a form of explicit data replay to preserve the model’s proficiency in these high-resource languages, as these are the most commonly utilized foreign languages in Hungary. The code component was included to maintain the model’s strong reasoning and structured-text processing abilities, which are known to be beneficial for a wide range of downstream analytical tasks (ma2024at).
Table 1: Composition of the Racka Pretraining Corpus after subset weighting and tokenization of the Hungarian-adapted tokenizer.
Language BPE Tokens (Billions) Ratio Documents (Millions)
Hungarian $\sim$70 44% 70.56
English $\sim$38 24% 26.12
German $\sim$34 21% 24.06
Code $\sim$18 11% 10.17
Total $\sim$160 100% 130.91
3.2 Data Sources
The corpus was assembled from a combination of large-scale web crawls and high-quality curated datasets to ensure both breadth and quality. The primary source for web data across all languages was the Common Crawl repository, with all data sourced from crawls conducted prior to our cutoff date of October 2024. (commoncrawl2025)
For English, we sample approximately 3838B tokens sourced from non-web text-based subsets of The Pile (gao2020pile800gbdatasetdiverse), and the heavily filtered fineweb corpus (penedo2024finewebdatasetsdecantingweb). We sample approximately 3434B tokens of German text from the similarly pre-filtered occiglot-fineweb (occiglotfineweb2024) and 1818B tokens of code from The Stackv2 (lozhkov2024starcoder2stackv2), slightly oversampling structured text formats and markups to better represent them compared to code files. 7070B Hungarian tokens were collected from various sources including Common Crawl, academic repositories, court rulings, movie subtitles, Wikipedia and digital news articles. Following Nemeskey:2020, the web and news subcorpora were heavily filtered and deduplicated on both the URL and document-level. Paragraphs that appeared frequently across the respective subcorpora were also discarded. The aggregated per-language corpus sizes are reported in Table 1; the composition of the Hungarian corpus is detailed in Appendix B. Note that the fertility of our adapted tokenizer is smaller for Hungarian, thus we produce less tokens in Hungarian compared to the other subsets.
In case of Hungarian web data, we apply standard boilerplate filtering and a strict paragraph-level 13-gram (whitespace token) deduplication resulting in our final dataset size. To ensure that the prevalence of scarce but high quality data is higher in the training dataset, we oversample the news subset (weight $2.0$), the Wikipedia subset (weight $3.0$), Hungarian subtitles (weight $2.0$) and academic repositories (weight $1.5$). We drop those documents that are less than $500$ characters long but do not apply other length-based filtering.
For measuring training efficiency we hold out 10-10 thousand element validation and test sets of independent documents. Both sets are sampled in a language-stratified manner.
4 Methodology
4.1 Backbone Model Architecture
The foundation for our work is the Qwen-3 4B model, an open-source transformer-based LLM developed by Alibaba Cloud accessed through Hugging Face [^1] https://huggingface.co/Qwen/Qwen3-4B. We selected this model as our backbone due to its strong baseline performance, favorable open-source license (Apache 2.0), and an architecture well-suited for our objectives. The Qwen-3 4B model features $4$ billion total parameters ($3.6$ billion non-embedding) distributed across $36$ transformer layers (yang2025qwen3technicalreport).
Key architectural specifications include the use of Grouped Query Attention (GQA) for efficient inference, the SwiGLU activation function for improved performance, and RMSNorm for stable training. (yang2025qwen3technicalreport) The model supports a native context length of 32,768 tokens, making it capable of processing long documents without truncation, which could be extended via interpolation of the positional encoding. (su2023roformerenhancedtransformerrotary) We specifically chose the reasoning and instruction-tuned variant of the model, hoping that its inherent capabilities would provide a stronger starting point and ultimately benefit in downstream performance after our continual pretraining phase. To our knowledge, this makes Racka the first reasoning model to be adapted to Hungarian.
4.2 Tokenizer Adaptation for Hungarian
A key hypothesis of our work is that the default tokenizer of a multilingual model is suboptimal for a morphologically rich, agglutinative language like Hungarian. Such tokenizers, trained on a distribution dominated by English, tend to fragment Hungarian words into numerous, less meaningful subword units, leading to longer sequence lengths, increased computational cost, and potentially poorer model performance (csaki2023efficientlyadaptingpretrainedlanguage). Several previous models have also experimented with tokenizer adaptation. To address this issue, we developed a new, adapted tokenizer for Racka, slightly modifying previously used adaptation algorithms. The process involved several steps:
  1. 1.
    We first trained a new Byte-Pair Encoding (BPE) tokenizer from scratch, using only a small portion of our Hungarian training corpus. This produced a vocabulary optimized for the statistical properties of Hungarian.
  2. 2.
    We then merged this new Hungarian-specific vocabulary with the original vocabulary of the Qwen-3 tokenizer. Non-latin character-based multi-byte tokens were removed, and novel tokens were inserted resulting in an adapted vocabulary that retained all of the original latin alphabet tokens while adding 32 00032\ 000 new, Hungarian-optimized tokens learned from the Hungarian-only tokenizer. See Appendix A for details.
  3. 3.
    The model’s architecture was updated to accommodate this change: specifically, the token embedding matrix and the language-model head (the final output layer) were altered to match the new tokenizer. To provide a strong initialization for the newly added tokens in the embedding matrix, we employed the VIPI (Vocabulary Initialization with Partial Inheritance) vocabulary-transfer technique (mosin2022finetuning). For each new token, we used the original tokenizer to break it down into its constituent subwords. The new token’s initial embedding vector was then calculated as the average of the embedding vectors of these constituent subwords from the original, pretrained embedding matrix. This approach provides a semantically meaningful starting point for the new tokens, facilitating faster convergence during training.
4.3 Packed Continual Pretraining with LoRA
To enable parameter-efficient continual pretraining, we employed Low-Rank Adaptation (LoRA) adapters. LoRA modules were inserted into all linear layers of the Qwen3-4B backbone. The original backbone weights remained frozen throughout training. The embedding and language modeling head layers are tied via transposed weight replication as in the Qwen3-4B base.
Our choice of LoRA hyperparameters was guided by recent best practices and empirical validation. Specifically, we set the LoRA rank ($r$) to $64$ and the scaling factor ($\alpha$) to $128$, following the heuristic of setting $\alpha$ to approximately twice the rank to ensure sufficient adaptation capacity. A dropout of $0.1$ was applied to the LoRA adapters to improve generalization. With this setup we train 0.520.52B parameters accounting for $12.5\%$ of the base model size. We used separate learning rates for LoRA and non-LoRA parameters: $1\times 10^{-4}$ for LoRA and $5\times 10^{-5}$ for non-LoRA parameters, both with a weight decay of $0.005$. Optimization was performed using AdamW. The learning rate schedule included a linear warmup phase covering $1\%$ of total training steps, followed by linear decay to zero which is shown to be beneficial (bergsma2025straightzerolinearlydecaying). We did not employ a specialized tokenizer healing phase.
To maximize training efficiency, we used input sequence packing (ding2024fewertruncationsimprovelanguage), merging shorter sequences into a single context window of up to 4096 tokens. Both the beginning-of-sequence and end-of-sequence delimiters were set to <|endoftext|>, as defined by the Qwen tokenizer. Packing was performed using a linear scaling algorithm prior to training. For computational efficiency, we used a standard full causal attention mask (as in (deepseekai2025deepseekv3technicalreport)), rather than a multi-segment lower-triangular mask, which would otherwise require quadratic memory and computational overhead (poedator4dmasks2024).
4.4 Training Infrastructure and Parallelization Strategy
Training was implemented in Python 3.12 using PyTorch 2.7.1 and CUDA 12.6, with the SDPA backend for efficient attention computation on NVIDIA A100 GPUs. All experiments were conducted using the HuggingFace Transformers library version 4.52.3, Accelerate 1.7.0, and PEFT 0.15.2.
[…]
Given the 4B parameter size of our model, which fits within the 40 GB memory of a single A100 GPU using gradient checkpointing, we empirically compared Fully Sharded Data Parallel (FSDP) and Distributed Data Parallel (DDP) strategies. While FSDP reduces per-GPU memory usage by sharding model states, it incurs frequent, latency-sensitive all-gather operations, which proved to be a bottleneck on Komondor’s interconnect. In contrast, DDP replicates the model on each GPU and synchronizes gradients via a single, bulk all-reduce after each backward pass, resulting in less frequent communication.
We conducted performance measurements on a distributed setup consisting of two nodes, each equipped with four GPUs. In DDP configuration, the training throughput reached 2.92.93.03.0 seconds per iteration. In contrast, FSDP configuration delivered substantially lower performance, reaching 5.05.05.15.1 seconds per iteration. Mixed-precision training with the standard FSDP implementation is not possible, training with full precision resulted in OOM for our minimal training context length of 44k tokens. As DDP yielded superior scaling and GPU utilization, so we adopted it for all experiments as a hardware-aware choice for parallelization on the available infrastructure.
Training was performed on $64$ A100 GPUs ($16$ nodes) with a per-device batch size of $2$ throughout $287$ hours resulting in a total GPU time of $2.1$ years. During training we experienced $1$ unexpected (hardware failure) and $2$ scheduled stops which resulted in no significant loss of compute with checkpoints in place. The estimated carbon emission for the training is $1290\text{ kg CO}_{2}\text{ eq.}$ which is a below average emission due to the environmentally friendly design of the HPC utilized.
4.5 Evaluation Protocol
Our evaluation protocol was designed to assess the model from two perspectives: its intrinsic ability to model language distributions and its practical capabilities on different Hungarian language model benchmarks. In this section we will introduce the intrinsic evalution and the mentioned benchmarks one-by-one.
Intrinsic Evaluation: We use perplexity (PPL) as our primary intrinsic metric. Perplexity measures the model’s uncertainty in predicting the next token in a sequence; a lower perplexity score indicates a better fit to the data distribution. (10.1121/1.2016299) We calculated perplexity on the held-out validation set containing a mix of Hungarian, English, German and code to monitor learning progress and assess for catastrophic forgetting.
LM-Eval-Harness benchmark: To measure the model’s practical utility without task-specific training, we conducted few-shot evaluations on a selection of Hungarian language understanding tasks (eval-harness). For this we utilized the lm-eval-harness with Hungarian translated benchmarks. We used the already implemented benchmarks ARC_HU, MMLU_HU, Hellaswag_HU and TruthfulQA_HU from lm-eval-harness, which is the work of (okapi). Additionally, we manually added the hungarian translated version of GSM8K (cobbe2021gsm8k), which is a mathematical reasoning dataset and was transletd by (thellmann2024crosslingual), to check the model reasoning capabilities on mathemical problems.
We ran all benchmarks separately, using the default config parameters defined in the original lm-eval-harness repository [^2] https://github.com/EleutherAI/lm-evaluation-harness. Table 3 shows our results.
HULU Benchmark The HULU benchmark offers a comprehensive set of supervised tasks covering a wide range of linguistic tasks, including acceptability, causal reasoning, sentiment classification, pronoun resolution, factual commitment, and reading comprehension. For the evaluation, we separately fine-tuned our model on each HULU subtask using a forked version of the benchmark repository [^3] https://github.com/nytud/HuLU. Fine-tuning was performed using full finetune and Low-Rank Adaptation (LoRA) matrices with the following key parameters: lora_rank=64lora\_rank=64, lora_alpha=128lora\_alpha=128, lora_dropout=0.1lora\_dropout=0.1, and batch_size=2batch\_size=2.
OpenHuEval benchmark To further evaluate the ability of the model to perform Hungarian reading comprehension and generation tasks, we conducted fine-tuned downstream evaluations using the OpenHuEval benchmark suite (openhueval). This collection focuses on cloze-style question answering and comprehension tasks designed to capture various aspects of contextual understanding, reasoning, and lexical inference in Hungarian. OpenHuEval comprises five datasets HuWildBench (real Hungarian user queries), HuSimpleQA (fact-based Hungarian questions), HuProverbRea (reasoning with Hungarian proverbs), HuMatchingFIB (option-based fill-in-the-blank), and HuStandardFIB (free-form fill-in-the-blank). The evaluation was made using the original repository of the benchmark [^4] https://github.com/opendatalab/OpenHuEval.
For comparison we evaluated the original Qwen3-4B, the Qwen3-4B-Base and the PULI-LlumiX-Llama-3.1 models on all benchmarks to check the development of the model after the training and to compare it with another hungarian LLM. The evaluation of the models were on a single A100 GPU for all benchmarks.
5 Results and Analysis
5.1 Tokenizer Performance
The primary goal of our tokenizer adaptation was to improve the encoding efficiency for Hungarian without degrading performance on the high-resource languages present in the base model. We evaluated this using subword fertility, which is the average number of tokens required to represent a single whitespace delimited word.
The results measured on the validation documents, presented in Table 2, confirm the success of our adaptation strategy. For Hungarian, the Racka tokenizer demonstrates a substantial improvement over the original Qwen-3 tokenizer. Subword fertility was reduced by over $46\%$, from $3.13$ down to $1.66$. This means that, on average, the same Hungarian text can be represented with almost half of the original Qwen tokens, leading to direct improvements in inference efficiency and latency. Crucially, this gain did not come at the same expense of performance on other languages with slight increase for German and a less than $24\%$ fertility increase for English. This indicates that our vocabulary extension successfully preserved the efficiency for high-resource languages.
Table 2: Tokenizer performance on the validation set: Subword fertility (lower is better) and relative change from Qwen-3 4B.
Language Qwen-3 4B Racka-4B Change (%)
Hungarian 3.1269 1.6584 -46.96
English 1.5705 1.9387 23.44
German 2.0502 2.3090 12.62
5.2 Continual Pretraining Results
Following the successful tokenizer adaptation, we successfully conducted the final continual pretraining phase on the full 160B token dataset. The training ran for 326 357326\,357 steps on 64 GPUs in parallel, with a per-device batch size of $2$ and gradient accumulation set to $4$ resulting in an effective logical batch size of $512$ sequences with a context length of $4096$ for each model update. The training loss curve is depicted in 1 Figure. We see that even though the training loss begins to flatten after 50 00050\,000 steps, the validation perplexity continues to decrease, indicating that overfitting is not happening.
Figure 1: Average training Categorical-Crossentropy loss logged every $200$ steps (left) and validation perplexity measured every 12 00012\ 000 steps on a language-stratified held out set of documents (right).
We evaluated the resulting model, Racka, in a two-shot setting on selected tasks from the llm-eval-harness Hungarian benchmark and compared its performance against the original Qwen-3 4B base model. The results are presented in Table 3. The preliminary outcomes, as anticipated in our abstract, are modest, but we see great potential in our models future training and downstream applicability. We could also see a drop from the original Qwen3-4B baseline,in the case of ARC (hu), where the original model attained a score of $32.02$.
Table 3: LM-Eval-Harness results for evaluated models (best results in bold)
Dataset (Metric) Qwen3-4B Racka-4B Qwen3-4B-base
Arc_hu (Acc) $0.3202$ $0.3005$ $0.3792$
Arc_hu (Acc_norm) $0.3844$ $0.3450$ $0.4169$
Hellaswag_hu (Acc) $0.3369$ $0.3407$ $0.3610$
Hellaswag_hu (Acc_norm) $0.4095$ $0.4078$ $0.4557$
MMLU_hu (Acc) $0.5427$ $0.5089$ $0.5965$
TruthfulQA_hu_mc1 (Acc) $0.3177$ $0.3047$ $0.3281$
TruthfulQA_hu_mc2 (Acc) $0.5060$ $0.5040$ $0.5045$
GSM8K_hu (Strict-match) $0.6330$ $0.5299$ $0.6398$
GSM8K_hu (Flexible extract) $0.6285$ $0.5329$ $0.6421$
The HuLU benchmark paints a more positive picture (see Table 4). Racka seems to be behind the base model(s) on HuRTE and HuCommitmentBank, but outperforms them on the other four tasks. On HuWNLI and HuCommitmentBank, it even surpasses the current state-of-the-art 8B model and achieves the best result out of all tested models on the former [^5] All numbers are measured on the validation set..
Table 4: Result of the mentioned 3 models on the datasets of the HULU benchmark, values averaged over multiple runs. Best result per benchmark is highlighted with bold.
Dataset Qwen3-4B Racka-4B Qwen3-4B-Base PULI-LlumiX-Llama-3.1 8B
ACC MCC F1 ACC MCC F1 ACC MCC F1 ACC MCC F1
HuCOLA $0.8109$ $0.3482$ $0.7840$ $0.8624$ $0.5657$ $0.8563$ $0.8254$ $0.4044$ $0.8027$ $0.8989$ $0.6920$ $0.8969$
HuCOPA $0.5589$ $0.1181$ $0.5584$ $0.7990$ $0.5998$ $0.7988$ $0.5845$ $0.1705$ $0.5837$ $0.9359$ $0.8720$ $0.9359$
HuSST $0.7517$ $0.5022$ $0.7433$ $0.7603$ $0.5137$ $0.7511$ $0.7539$ $0.5082$ $0.7513$ $0.7804$ $0.5598$ $0.7698$
HuRTE $0.9078$ $0.8142$ $0.9078$ $0.8790$ $0.7553$ $0.8790$ $0.8872$ $0.7719$ $0.8872$ $0.8979$ $0.7936$ 0.897 690.897\,69
HuWNLI $0.5033$ 0.098-0.098 $0.3862$ $0.5666$ $0.1031$ $0.4548$ $0.5366$ 0.060-0.060 $0.4069$ $0.38$ 0.2815-0.2815 $0.3668$
HuCommitmentBank $0.7378$ $0.6078$ $0.7316$ $0.6388$ $0.4741$ $0.6373$ $0.6291$ $0.4733$ $0.6112$ $0.4854$ $0.2742$ $0.4594$
Overall Performance $0.711$ $0.382$ $0.685$ $0.751$ $0.502$ $0.7295$ $0.702$ $0.378$ $0.673$ $0.729$ $0.485$ $0.721$
Finally, Table 5 presents the results of evaluating Racka on generation tasks. Beyond the base model and our own, we also compare against the models reported in the original OpenHuEval paper that are close in size. The results show that Racka delivers competitive performance relative to models twice its size and to the corresponding base models.
Table 5: Results on the datasets of the OpenHuEval Benchmark. Best result per benchmark is highlighted with bold.
†: After careful evaluations, we have encountered a problem with the HuWildBench prompting methods and tokenizer compatibility. We will include the revised results in the final paper as well.
Model HuWildBench HuSimpleQA HuProverbRea HuMatchingFIB HuStandardFIB Overall
WBScore Acc Acc (OE) Acc (2CQ) B Acc, Q Acc B Acc Q Acc
Llama-3.1-Instruct-8B $53.62$ $15.2$ $63.35$ $73.48$ $5.74$ $0.72$ $16.64$ $1.08$ $28.72$
Qwen2.5-Instruct-7B $42.01$ $5.22$ $50.48$ $67.05$ $31.88$ $1.08$ $7.43$ 0 $25.64$
Qwen3-4B t.b.d.† $7.3$ $62.47$ $69.25$ $8.1$ $4.14$ $4.68$ $2.15$ t.b.d.†
Qwen3-4B-Base t.b.d.† $5.9$ $41.15$ 0 $42.3$ $5.58$ 0 0 t.b.d.†
Racka-4B t.b.d.† $7.6$ $58.41$ $77.53$ $22.87$ $2.52$ $18.02$ $1.08$ t.b.d.†
6 Conclusion and Future Directions
This work introduced Racka, a resource-efficient, continually pretrained language model for Hungarian, using a practical methodology specifically designed for deployment on national HPC infrastructure. Through targeted tokenizer adaptation by expanding and initializing the vocabulary to better capture the morphological complexity of Hungarian, we achieved a substantial reduction in subword fertility, improving both computational efficiency and the representational capacity of the model. The construction of a 160B-token multilingual corpus, with a carefully balanced data mixture, further enabled robust language adaptation while reducing catastrophic forgetting of high-resource languages.
The evaluation results partially confirm the effectiveness of our tokenizer adaptation and training strategy. Racka outperforms its base model on some tasks and lags behind it on others. The same relationship can be seen between the reasoning Qwen3-4B and the foundational Qwen3-4B-base models, resulting a complicated picture. Nevertheless, Racka seems to be competitive on most of the benchmarks, sometimes with models twice its size.
Future work includes systematic hyperparameter optimization and supervised fine-tuning for downstream applications as well as preference tuning. We intend to openly publish our model and tokenizer, to support community-driven research and contribute to the advancement of digital sovereignty of Hungary.
Acknowledgement
We would like to thank Levente Szabados for the name idea and initial informal discussions.
This research was supported by the EKÖP-24 University Excellence Scholarship Program of the Ministry for Culture and Innovation, funded by the National Research, Development and Innovation Fund.
The authors acknowledge the support of the National Laboratory for Digital Heritage. Project no. 2022-2.1.1-NL-2022-00009 has been implemented with the support provided by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation Fund, financed under the 2022-2.1.1-NL funding scheme.
We acknowledge the Digital Government Development and Project Management Ltd. for awarding us access to the Komondor HPC facility based in Hungary.
Appendix
A Tokenizer Adaptation Algorithm
Algorithm 1 Extended Tokenizer Construction for Hungarian Adaptation
1:Original vocabulary $\mathcal{V}$ with size $m$; Hungarian corpus $\mathcal{C}$; target increment n=32,000n=32{,}000; target size $m^{\star}=m+n$
2:Extended vocabulary $\mathcal{V}_{\mathrm{ext}}$ of size $m^{\star}$
3:Train a BPE tokenizer of size $m^{\star}$ on $\mathcal{C}$ to obtain the ordered vocabulary $\mathcal{V}^{\prime}=(p_{0},p_{1},\dots,p_{m^{\star}-1})$.
4:Let $k$ be the index of the $n$-th token $p_{k}$ in $\mathcal{V}^{\prime}$ such that $p_{k}\notin\mathcal{V}$.
5:if $k=m^{\star}-1$ then
6:  return $\mathcal{V}^{\prime}$
7:else
8:  $\mathcal{V}^{\prime\prime}\leftarrow$ all tokens in $\mathcal{V}$ not in $\mathcal{V}^{\prime}[:k+1]$, ordered by $\mathcal{V}$’s original merge order
9:  return $\mathcal{V}^{\prime}[:k+1]$ concatenated with $\mathcal{V}^{\prime\prime}$
10:end if
B Composition of the Hungarian Corpus
The table below shows the composition of the Hungarian corpus with the corresponding number of documents and BPE tokens. For Common Crawl and News, the numbers shown are after filtering and deduplication. The former represent about 30% of the pre-filtered data.
Dataset Subset Documents (thousands) Tokens (millions) Percentage
Common Crawl .hu 52 775.31252\,775.312 40 621.051 62340\,621.051\,623 72.70872.708%
.ro $552.187$ 397.695 476 057397.695\,476\,057
.sk $446.774$ 224.604 267 996224.604\,267\,996
.com $2984.912$ 2641.126 134 562641.126\,134\,56
Repo EPA $300.790$ 2672.379 987 722672.379\,987\,72 19.478 97219.478\,972%
Books $32.830$ 1338.894 340 931338.894\,340\,93
OJS $27.702$ 96.362 846 119 296.362\,846\,119\,2
OAI $349.398$ 7649.321 370 747649.321\,370\,74
News $5938.789$ 2723.869 835 032723.869\,835\,03 4.54.5%
Court $198.296$ 1110.304 357 31110.304\,357\,3 3.33.3%
HuParl $1.707$ 140.560 154 724140.560\,154\,724
OpenSubtitles $88.519$ 508.851 908 561508.851\,908\,561
Wikipedia $158.457$ 232.159 235 234232.159\,235\,234
Sum 63 855.67363\,855.673 60 357.181 537 960\,357.181\,537\,9 100100%